If cos A = - 5/13 with A in QIII, find cot 2A.
Answer by lwsshak3(11628)

If cos A = - 5/13 with A in QIII, find cot 2A.
***
cos A=-5/13
sin A=-12/13
cos 2A=cos^2A-sin^2A=25/169-144/169=-119/169
sin 2A=2sin A cos A=2(-12/13)(-5/13)=120/169
cot 2A=cos 2A/sin 2A=-119/120
..
Check:
cos A=-5/13 (In Q3)
A=247.38˚
2A=494.76˚
tan 2A=tan(494.76)=-1.0084=1/cot 2A
cot A≈-0.9916 (w/calculator)
exact value=-119/120≈-0.9916